1.The management of the Lackawanna Steel Company has estimated the following elasticities for a special type of steel they sell: Own price elasticity of demand = -2, Income elasticity of demand = 1 and Cross price elasticity of demand with respect to aluminum = 1.5. The management of the firm has forecasted that next year income in the relevant market segment will increase by 3 % and that the price of Aluminum will fail by 2%. The firm is planning to increase the price of its steel by 4 %. If the sales of the firm this year are 2000 tons of this special type of steel, how many tons of steel can the firm expect to sell next year (after the proposed price increase)? (Hint: Assume that the income and price effects are independent and additive.)
Answer #1

Ans. We know that , Price Elasticity of Demand = % change in demand ÷ % change in price

% change in quantity = Price Elasticity × % change in Price

% change in quantity = -2 × 4%

% decrease in quantity = - 8%

2000 × 8 / 100 = 160 ; 2000 - 160 = 1840

Firm can expect to sell 1840 tons of steel next year.

Best of Luck !! !!